OpenStudy (raphibomb):

erica hiked up a mountain trial at 3 km/h and returned at 4 km/h. the entire trip took 5 hours and 10 minuites, including the half-hour she spent at the top. how long was the trail how does this work

OpenStudy (3mar):

If I may help?

OpenStudy (raphibomb):

?

OpenStudy (raphibomb):

can u? please...

OpenStudy (3mar):

Thank you.

OpenStudy (raphibomb):

mp

OpenStudy (3mar):

entire trip took 5 hours and 10 minuites, what can you conclude from that?

OpenStudy (raphibomb):

that it was 310 min?

OpenStudy (3mar):

Great!

OpenStudy (raphibomb):

yay!

OpenStudy (3mar):

"including the half-hour she spent at the top", what is the net time for just go and back?

OpenStudy (raphibomb):

280 min

OpenStudy (raphibomb):

oh

OpenStudy (raphibomb):

ye

OpenStudy (3mar):

Excellent!

OpenStudy (3mar):

what is the speed for going?

OpenStudy (raphibomb):

3 km/pm

OpenStudy (3mar):

3 km/hr you mean, ok? what in km/min?

OpenStudy (raphibomb):

180km/pm?

OpenStudy (raphibomb):

ye i meant hours but thought it would be easier in minutes...

OpenStudy (3mar):

\[3km/hr=\frac{ 3 }{ 60 }km/\min\]

OpenStudy (raphibomb):

ok.

OpenStudy (3mar):

what for back?

OpenStudy (raphibomb):

4/60 km/min

OpenStudy (3mar):

and we know that \[speed=\frac{ distance }{ time }\] so that \[diatance=speed×time\] ok?

OpenStudy (raphibomb):

ok

OpenStudy (3mar):

and she go and come back the same distance twice ok?

OpenStudy (raphibomb):

ok

OpenStudy (3mar):



OpenStudy (3mar):

so now we can equate the (speed×time) for the go one to the back one can you share your work?

OpenStudy (raphibomb):

so it's \[3=x/280\]

OpenStudy (raphibomb):

or 3+x/280=4+x/280

OpenStudy (3mar):

No man 280 is for both together not for one trip

OpenStudy (raphibomb):

oh ok...

OpenStudy (3mar):

I mean\[v_1*t_1=v_2*t_2\] and we have v_1 and v_2 can you show me?

OpenStudy (raphibomb):

whats v and t

OpenStudy (3mar):

v stands for speed and t for time and x for distance

OpenStudy (3mar):

Can you figure out the relation between t_1 and t_2?

OpenStudy (3mar):

for km/min: \[\frac{ 3 }{ 60 }*t_1=\frac{ 4 }{ 60 }*t_2\] ok?

OpenStudy (raphibomb):

ok, great

OpenStudy (raphibomb):

so now do we need to find out the time?

OpenStudy (3mar):

1 min plz

OpenStudy (raphibomb):

ok

OpenStudy (3mar):

Sorry for late.

OpenStudy (3mar):

So as we get the final equation, we can substitute with it into the first one!

OpenStudy (3mar):

\[\frac{ 3 }{ 60 }*t_1=\frac{ 4 }{ 60 }*t_2........3t_1=4t_2..............\huge t_1=\frac{ 4 }{ 3 }t_2\]

OpenStudy (3mar):

@raphibomb Can you proceed?

OpenStudy (3mar):

\[\huge t_1+t_2=280\min\]

OpenStudy (3mar):

\[\huge t_1=\frac{ 4 }{ 3 }t_2\] \[\huge t_1+t_2=280\min\] \[\frac{ 4 }{ 3 }t_2+t_2=280\] \[\frac{ 7 }{ 3 }t_2=280\] \[t_2=120\min\] subsitute in any equation, we get t_1: \[t_1=\frac{ 4 }{ 3 }t_2=\frac{ 4 }{ 3 }*120=160\min\] For check: to go =120 min to come back = 160 min she stayed here for= 30 min add theses periods up, you got 310 as we calculated firstly

OpenStudy (3mar):

I am happy to hear that.

OpenStudy (3mar):

@raphibomb If you have any difficulties, kindly tell me!

OpenStudy (3mar):

Salam!